OpenStudy (anonymous):

N2 is expelled from a vessel of 5LCapacity by heating it from 27 C to 57C the fraction of N2 gas explled under condition that the pressure is Constant at 2 atm?

OpenStudy (ghazi):

do yo have to calculate amount of gas expelled ?

OpenStudy (anonymous):

Percentage...

OpenStudy (ghazi):

use \[PV=nR \Delta T\] find n and then assume that initially you had one mole of it and you can find the percentage

OpenStudy (ghazi):

your n will be around 0.04 mole

OpenStudy (anonymous):

It wont match the answer shuld be 10%

OpenStudy (ghazi):

hmm

OpenStudy (ghazi):

what did you use at the place of R

OpenStudy (anonymous):

0.0821

OpenStudy (ghazi):

i think R= 8.314 should be used

OpenStudy (ghazi):

but i am sure @Kryten will give the answer :) so wait

OpenStudy (anonymous):

ok so start with this: pV=nRT since there is no change in amount: n1=n2 (pV/RT)1=(pV/RT)2 now since R is an constant you can cross it out and you get (pV/T)1 =(pV/T)2 now since there is no change in pressure or in other words it is constant you can cross it out V1/T1 = V2/T2 now all you need to do is to use that equation to calculate how much did N2 expand and then: V2 = ??? Vx= V2-V1 and V(%) = Vx/V1

OpenStudy (anonymous):

Thxx.....).....